OpenStudy (anonymous):

fiind the center and radius of the circle whose equation is given: x^2+y^2+2x-3=0

OpenStudy (mertsj):

\[x^2+2x +1 +y^2=3+1\] \[(x+1)^2+(y-0)^2=4\]

OpenStudy (mertsj):

center (-1,0) r =2